Nickel carbonyl has the formula Ni(CO)x. The relative molecular formula mass nickel carbonyl is 171. Calculate the value of x.
1
Expert's answer
2020-07-14T15:19:25-0400

Molar mass of nickel is 59 and molar mass of CO group is 28 =>171-59=112/28=4=> x=4

Ni(CO)4